OpenStudy (anonymous):

Find the intercept: 6x-2y=-24

OpenStudy (anonymous):

first solve for y.

OpenStudy (anonymous):

is it 12?

OpenStudy (anonymous):

6x-2y=-24 -2y = 24-6x y = 3x +(-12) the y intercept is when x = 0. what is y

OpenStudy (anonymous):

x=4 & y=-12 ?

OpenStudy (anonymous):

y intercept is -12

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

you're welcome